Circular No. 6821

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1998 B1 (SHOEMAKER-LEVY 8) C. W. Hergenrother has recovered comet P/1992 G2 (= 1992f = 1992

XV) on images obtained with the Smithsonian Astrophysical Observatory’s 1.2-m reflector at Mt. Hopkins on Jan. 22. He confirmed the recovery on images obtained with the Lunar and Planetary Laboratory’s 1.5-m reflector at the Catalina Station on Jan. 28, when there was a moderately diffuse 5′′ coma but no tail on a 900-s exposure.

1998 UT  α_{2000}  δ_{2000}  m_{1}
Jan. 22.23128  7^{h}28^{m}13.^{s}12  +14^{o}00′30′′.7  21.8
22.23514  7 28 13.05  +14 00 30.7  21.7
22.23898  7 28 12.92  +14 00 30.5  22.0
28.34848  7 24 02.92  +14 11 32.2
28.37097  7 24 02.00  +14 11 35.3  22.6

Hergenrother’s measurements above show that the prediction by the undersigned on MPC 27082 (ephemeris on MPC 30702) requires correction by ΔT = +0.03 day. The following improved orbital elements represent 76 observations 1992–1998 with mean residual 0′′.8:

Epoch = 1999 Dec. 8.0 TT

T = 1999 Dec. 10.5994 TT  ω = 22.^{o}5940
e = 0.289451  Ω = 213.3174 }2000.0
q = 2.721257 AU  i = 6.0505

a = 3.829792 AU n^{o} = 0.1315047 P = 7.495 years

COMETS 55P/TEMPEL-TUTTLE AND 103P/HARTLEY 2 M. Fomenkova, J. Sarmecanic and M. Wang, University of California

at San Diego, report the following infrared magnitudes (uncertainties 0.20– 0.25 mag), obtained on Jan. 24 with the 1.5-m telescope on Mt. Lemmon (+ UCSD mid-infrared camera + 20′′ circular aperture): comet 55P, [8.7 μm] = 2.7, [10.3 μm] = 1.5, [11.7 μm] = 1.5, [12.5 μm] = 1.3; comet 103P, [10.3 μm] = 1.6, [11.7 μm] = 1.3, [12.5 μm] = 0.9. The magnitudes of 55P indicate a possible silicate excess of 0.4 mag and yield a color temperature of 300 ± 15 K, which is 15 percent higher than the blackbody temperature of 263 K at r = 1.13 AU. There was no evidence of excess silicate emission in the case of 103P.

1998 February 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden